# Integrating Deep-Learning-Based Magnetic Model and Non-Collinear Spin-Constrained Method: Methodology, Implementation and Application

### v2

@@ II.2 Projection Methods
$w_{I}(\mathbf{r})$ is the weighting function and partition function $P_{I}(\mathbf{r})$ approaches to 1 near $I$-th atom and 0 otherwise. A famous example of a more sophisticated partitioning strategy is called Atom-In-Molecule (AIM), or well-known Bader charge analysis Bader and Molecules (1990), which defines atomic regions by zero-flux surfaces in the gradient of the electron density, Voronoi tessellation Fonseca Guerra et al. (2004) partitions space into polyhedral cells around each atom based on proximity. The Bader charge analysis provides a mathematically rigorous definition of atomic boundaries based on electron density, yielding charges that are invariant to the choice of basis set and more reflective of the true electronic structure. However, this method would fail in cases where there are zero or more than one atoms appear in one Voronoi cell, and lack of decomposition of the orbits.
Mulliken Mulliken (1955) and Löwdin Löwdin (1970) charge analysis methods are well-established for partitioning electron density within a molecule to assign atomic charges. The formula for calculating the atomic magnetization from Mulliken charge of $I$-th atom is given by
@@
The subspace projection method requires a rational construction of projection operators in which the projection function should be of physical meaning. For example in projected augmented wave (PAW) formulism Blöchl (1994), the PAW projector $\{|p_{I\mu}\rangle\}$ which holds the orthonormality with the pseudo partial wave that is connected with the all-electron one $\{|\phi_{\mu}\rangle\}$ is a natural choice Ma and Dudarev (2015); Hegde et al. (2020a). Atomic magnetic moments is defined as
@@
where the radial functions $\alpha(r)$ are constructed to meet three criteria: (1) maximized efficiency for extracting valence electron information of interest, (2) normalization, and (3) smoothness at the boundary to avoid numerical error. Practically, we construct smooth modulation orbitals (SMOs) by truncating those $\zeta$ functions of ABACUS numerical atomic orbitals (NAOs) Lin et al. (2021); Chen et al. (2010) that collect majority of valence electron information over various reference systems, then smoothed by a function centered at the boundary
@@
where $r$ is the distance to the atom, $r_{m}$ is the artificially chosen modulation radius. Spreading parameter $\sigma$ is solved in an iterative way, that can maximize the overlap integral between the SMOs and NAOs under the constraint of normalization of the SMOs themselves
$\alpha(r)=\left.\frac{\chi(r)g(r;\sigma)}{\langle\chi(r)g(r;\sigma)|\chi(r)g(r;\sigma)\rangle^{1/2}}\right|_{\frac{\partial\langle\alpha|\chi\rangle}{\partial\sigma}=0},$ (23)
where $\chi(r)$ are the radial parts of $\zeta$ functions in NAOs. Using SMOs as projection orbitals allows for the representation of charge distributions across different angular momentum orbitals within an atom’s localized environment. Since the projection orbitals are designed to be strongly localized, inter-atomic overlap can be neglected. Furthermore, the orthogonality of projection orbitals with different angular momenta is preserved during modulation, resulting in a simplified form for the projection operator.
@@
is the plane wave basis representation of the overlap between $|\psi_{n\mathbf{k}}^{\sigma}\rangle$ and SMOs, $c_{n\mathbf{k}}^{\sigma}(\mathbf{G})$ and $\alpha_{Ilm}(\mathbf{G})$ are coefficients of plane wave expansion of $|\psi_{n\mathbf{k}}^{\sigma}\rangle$ and SMOs, respectively. One can refer to Ref. King-Smith et al., 1991 for the details of SMOs expansion.
II.3 Spin-Constrained DFT
Figure 1: Framework of variant Lagrange quantities spin-cDFT algorithm implemented in ABACUS for both PW and NAOs basis. The electron density is $\rho$ and the spin density vector is $\mathbf{m}$. The atomic magnetization is $\mathbf{M}$.
Utilizing the established Lagrange formalism Dederichs et al. (1984b); Cai et al. (2023), the challenge pertaining to the restriction of atomic magnetic moments is reformulated as an endeavor to ascertain the equilibrium points of the pertinent function. To this end, a Lagrange multiplier is introduced into the energy functional of the system:
@@ II.3 Spin-Constrained DFT
where $E_{\text{KS}}$ is the Kohn-Sham energy with charge density $\rho(\mathbf{r})$ and spin density $\mathbf{m}(\mathbf{r})$, $\bm{\lambda}_{I}$ is the Lagrange multiplier, which can be treated as the magnetic torque under this constraint. $\mathbf{M}_{I}$ and $\mathbf{M}_{I,\text{target}}$ are the atomic spin moment and the target atomic spin moment for atom $I$, respectively. The stationary point problem within the Lagrange formalism can be solved iteratively by minimizing $E_{c}$ with respect to $\bm{\lambda}_{I}$, $\rho(\mathbf{r})$ and $\mathbf{m}(\mathbf{r})$ Wu and Van Voorhis (2005).
II.3.1 Numerical Atomic Orbital Basis
The penalty Hamiltonian term in NAOs basis in real space can be derived from the Lagrange function as
@@ II.3.1 Numerical Atomic Orbital Basis
where the 2$\times$2 parameter matrix $f(I,\sigma\sigma^{\prime})$ from Pauli matrix is
@@
II.3.2 Plane Wave Basis
Given a large number of plane wave basis functions, directly constructing the full Hamiltonian matrix is generally impractical. Instead, iterative diagonalization methods are employed to solve the Kohn-Sham equations, which require only the computation of the action of the Hamiltonian operator on the wave function. As a result, within the plane-wave basis set, the formula for the Hamiltonian correction term at a specified $\mathbf{k}$-point is as follows
@@ II.3.2 Plane Wave Basis
where $\Omega$ is the cell volume, $\varepsilon_{\alpha\beta}$ is the lattice strain tensor, and detail of last derivative term is similar with nonlocal pseudopotential term in Ref. Nielsen and Martin, 1985.
II.3.3 Workflow
@@
We present a uniform flowchart of the specific implementation of spin-cDFT in Fig. 1 for both PW and NAO basis. As depicted in Fig. 1, this procedure necessitates an additional iteration within the self-consistent field (SCF) cycle of DFT to ascertain the fixed-point solution satisfying the spin constraints. The process involves solving the Kohn-Sham equations with a given set of Lagrange multipliers, which are then updated to minimize the deviation between the calculated and target magnetic moments. This minimization is typically achieved through a conjugate-gradient scheme or a similar optimization technique, ensuring that the magnetic moments converge to the targets within a predefined tolerance. The iterative loop continues, alternating between the solution of the Kohn-Sham equations and the optimization of the Lagrange multiplier, until the system reaches a self-consistent solution where the constraints are satisfied to a desired level of accuracy.
III Results and Discussions
III.1 Computational Details
Figure 2: Radial functions (RFs) of the first-$\zeta$ numerical atomic orbital of Fe before and after modulation, where (a) $p$ orbitals and (b) $d$ orbitals display the radial functions of the original numerical atomic orbitals with a cutoff radius $r_{c}$ of 6.0 Bohr, as well as the modulated orbital with the modification radius $r_{m}$ ranging from 1.0 to 5.0 Bohr. “p-1.0” represents the $p$ orbital modulated by Eq. 22 with $r_{m}=1.0$ Bohr.

### v3

@@ 4.2 Projection Methods
$w_{I}(\mathbf{r})$ is the weighting function and partition function $P_{I}(\mathbf{r})$ approaches to 1 near $I$-th atom and 0 otherwise. A famous example of a more sophisticated partitioning strategy is called Atom-In-Molecule (AIM), or well-known Bader charge analysis [bader1990quantum], which defines atomic regions by zero-flux surfaces in the gradient of the electron density, Voronoi tessellation [fonseca2004voronoi] partitions space into polyhedral cells around each atom based on proximity. The Bader charge analysis provides a mathematically rigorous definition of atomic boundaries based on electron density, yielding charges that are invariant to the choice of basis set and more reflective of the true electronic structure. However, this method would fail in cases where there are zero or more than one atoms appear in one Voronoi cell, and lack of decomposition of the orbits.
Mulliken [mulliken1955electronic] and Löwdin [lowdin1970nonorthogonality] charge analysis methods are well-established for partitioning electron density within a molecule to assign atomic charges. The formula for calculating the atomic magnetization from Mulliken charge of $I$-th atom is given by
@@
The subspace projection method requires a rational construction of projection operators in which the projection function should be of physical meaning. For example in projected augmented wave (PAW) formulism [1994-paw], the PAW projector $\{|p_{I\mu}\rangle\}$ which holds the orthonormality with the pseudo partial wave that is connected with the all-electron one $\{|\phi_{\mu}\rangle\}$ is a natural choice [2015-ma, 2020-SPHInX]. Atomic magnetic moments is defined as
@@
where the radial functions $\alpha(r)$ are constructed to meet three criteria: (1) maximized efficiency for extracting valence electron information of interest, (2) normalization, and (3) smoothness at the boundary to avoid numerical error. Practically, we construct smooth modulation orbitals (SMOs) by truncating those $\zeta$ functions of ABACUS numerical atomic orbitals (NAOs) [lin2021strategy, Chen2010] that collect majority of valence electron information over various reference systems, then smoothed by a function centered at the boundary
@@
where $r$ is the distance to the atom, $r_{m}$ is the artificially chosen modulation radius. Spreading parameter $\sigma$ is solved in an iterative way, that can minimize the gradient term of spillage [lin_accuracy_2020] between the SMOs $|\alpha(\sigma)\rangle$ and NAOs $|\chi\rangle$ under the constraint of normalization of the SMOs themselves
$\min_{\sigma\in\mathbb{R}^{+}}\left\|\nabla|\chi\rangle-\nabla|\alpha\left(\sigma\right)\rangle\right\|^{2}\,\,\mathrm{s}.\mathrm{t}.\langle\alpha\left(\sigma\right)|\alpha\left(\sigma\right)\rangle=1,$ (23)
where
[…]
Using SMOs as projection orbitals allows for the representation of charge distributions across different angular momentum orbitals within an atom’s localized environment. Since the projection orbitals are designed to be strongly localized, inter-atomic overlap can be neglected. Furthermore, the orthogonality of projection orbitals with different angular momenta is preserved during modulation, resulting in a simplified form for the projection operator.
@@
is the plane wave basis representation of the overlap between $|\psi_{n\mathbf{k}}^{\sigma}\rangle$ and SMOs, $c_{n\mathbf{k}}^{\sigma}(\mathbf{G})$ and $\alpha_{Ilm}(\mathbf{G})$ are coefficients of plane wave expansion of $|\psi_{n\mathbf{k}}^{\sigma}\rangle$ and SMOs, respectively. One can refer to Ref. [king1991real] for the details of SMOs expansion.
4.3 Spin-Constrained DFT
Utilizing the established Lagrange formalism [dederichs1984ground, 2023-cai], the challenge pertaining to the restriction of atomic magnetic moments is reformulated as an endeavor to ascertain the equilibrium points of the pertinent function. To this end, a Lagrange multiplier is introduced into the energy functional of the system:
@@ 4.3 Spin-Constrained DFT
where $E_{\text{KS}}$ is the Kohn-Sham energy with charge density $\rho(\mathbf{r})$ and spin density $\mathbf{m}(\mathbf{r})$, $\bm{\lambda}_{I}$ is the Lagrange multiplier, which can be treated as the magnetic torque under this constraint. $\mathbf{M}_{I}$ and $\mathbf{M}_{I,\text{target}}$ are the atomic spin moment and the target atomic spin moment for atom $I$, respectively. The stationary point problem within the Lagrange formalism can be solved iteratively by minimizing $E_{c}$ with respect to $\bm{\lambda}_{I}$, $\rho(\mathbf{r})$ and $\mathbf{m}(\mathbf{r})$ [2005-wu].
This method can be extended in a straightforward manner to support a functionality that constrains only the direction of atomic spin moments. By rewriting the target atomic spin moment in Eq. 30 as $\mathbf{M}_{I,\text{target}}=M_{I,\text{target}}\mathbf{e}_{I,\text{target}},$ and iteratively updating the magnitude of the target spin moment $M_{I,\text{target}}=\mathbf{M}_{I}(m(r))\cdot\mathbf{e}_{I,target}$ during the inner-loop calculation of optimized Lagrange multipliers $\mathbf{\lambda}_{I}|_{\mathbf{\lambda}_{I}\cdot\mathbf{e}_{I,\text{target}}=0}$, upon the inner loop’s convergence, the following conditions are satisfied: $\mathbf{M}_{I}\big(\mathbf{m}(\mathbf{r})\big)\parallel\mathbf{e}_{I,\text{target}},$ and $\sum_{I}\lambda_{I}\cdot M_{I}=0.$ A typical example is the transition of a single atomic spin from ferromagnetic to antiferromagnetic testing, examining whether fixing the magnetic moment magnitude results in differences in energy and torque performance. The comparative studies between the full-constraint and direction-constraint method is perfomed in Fig. S6. It is worth noting that constraining only the spin direction-rather than both magnitude and direction-reduces the number of optimization degrees of freedom, typically cutting the required convergence steps in the inner loop by approximately half. Another noteworthy aspect is that the direction-only constraint approach circumvents errors arising from discrepancies in atomic magnetic moment definitions across different software packages, thereby enabling meaningful cross-software comparisons (see Fig. S14).
Another extension can be achieved by replacing all vector quantities in Eq. 30 with scalars, enabling the method to support spin-magnitude-constrained calculations under the LSDA. It is important to clarify that the two aforementioned extensions cannot be employed simultaneously, as the spin orientation in LSDA is restricted to two discrete directions (up and down), which cannot be successfully constrained through iterative gradient-based optimization.
First, we introduce the implementation within numerical atomic orbital basis. The penalty Hamiltonian term in NAOs basis in real space can be derived from the Lagrange function as
@@
where the $\times2$ parameter matrix $f(I,\sigma\sigma^{\prime})$ from Pauli matrix is
@@
For plane wave basis, directly constructing the full Hamiltonian matrix is typically infeasible due to the large number of basis functions involved. Instead, iterative diagonalization methods are employed to solve the Kohn-Sham equations, which require only the computation of the action of the Hamiltonian operator on the wave function. As a result, within the plane-wave basis set, the formula for the Hamiltonian correction term at a specified $\mathbf{k}$-point is as follows
@@
where $\Omega$ is the cell volume, $\varepsilon_{\alpha\beta}$ is the lattice strain tensor, and detail of last derivative term is similar with nonlocal pseudopotential term in Ref. [Nelson1985].
@@
We present a uniform flowchart of the specific implementation of spin-cDFT in Fig. 11 for both PW and NAO basis. As depicted in Fig. 11, this procedure necessitates an additional iteration within the self-consistent field (SCF) cycle of DFT to ascertain the fixed-point solution satisfying the spin constraints. The process involves solving the Kohn-Sham equations with a given set of Lagrange multipliers, which are then updated to minimize the deviation between the calculated and target magnetic moments. This minimization is typically achieved through a conjugate-gradient scheme or a similar optimization technique, ensuring that the magnetic moments converge to the targets within a predefined tolerance. The iterative loop continues, alternating between the solution of the Kohn-Sham equations and the optimization of the Lagrange multiplier, until the system reaches a self-consistent solution where the constraints are satisfied to a desired level of accuracy. Through comparative testing, we found that the energy calculation results from the adaptive algorithm implemented in this work demonstrate agreement with the penalty function methods employed in other software packages (see Fig. S14). Compared to the penalty function method, which exhibits nonlinear errors dependent on $\lambda$ across different excited states (Fig. S14), new method demonstrates consistent accuracy and is suitable for large-scale sampling.
Figure 11: The double-loop workflow of spin-cDFT. Framework of variant Lagrange quantities spin-cDFT algorithm implemented in ABACUS for both PW and NAOs basis. The electron density is $\rho$ and the spin density vector is $\mathbf{m}$. The atomic magnetization is $\mathbf{M}$.
We observed that in calculations of the Fe system, the optimization of lambda typically introduces an inner-loop iteration count on the order of 5. Since the inner loop also requires solving eigenvalue problems, this leads to an approximately fivefold increase in total computational time when using numerical atomic orbital basis sets (Fig. S15). For plane-wave basis set, due to its high dimensionality, the inner loop employs a single-step subspace solution of the Davidson diagonalization algorithm when the magnetic moment error is large, rather than the full Davidson solver used in the outer loop. This merit significantly improves the efficiency of magnetic moments in the inner loop while guaranteeing the correctness of final magnetic moment convergence. In tests, each step of the inner loop with PW basis sets consumes significantly less time than the outer loop, resulting in only a modest increase ($\sim 1/3$) in total computational time (see Fig. S7). Tests on systems of varying sizes (see Fig. S15) demonstrate that for both LCAO and PW basis implementations, the proportion of additional computational cost introduced by cDFT remains nearly constant as the system size increases. This indicates that the current cDFT method does not alter the scaling of the original SCF calculation; the extra cost only increases the prefactor of the total computational time, ensuring its scalability. Furthermore, the present method, due to its strict enforcement of magnetic moment constraints at each electronic step, reduces the degrees of freedom of the electronic state during the SCF process, leading to superior convergence speed compared to the penalty method (see Table. S6).
4.4 First-principles Calculations